3 dice are thrown simultaneously. The probability that 4 has appeared on 2 dice given that 5 has appeared on one dice

Grade:12

1 Answers

Arun
5 years ago
3 dthree dice are thrown simultaneously let A be the event that 4 appears on two dice and B be the event that 5 occures on exactly one die. We want to find conditional probability of A given B-
 
P(A/B) = P (A intersection B)/ P(B) = 3/216 / 75/216 = 1/25
 
Hope it helps